# Best of both worlds: Fusing hyperspectral data from two generations of spectro-imagers for X-ray astrophysics

### v1

@@ 2 Methodology
In this section, we present the methodology of our fusion algorithm for X-ray astrophysics. We go over the general problem at play in section 2.1. Section 2.2 goes over the rebinning operator needed when the data sets have different voxel sizes. Section 2.3 covers our choices of potential regularization terms. Section 2.4 goes over our algorithmic strategy to solve the inverse problem.
@@ 2.1 Forward Model
Let $X$ be the hyperspectral data from the instrument with better spatial resolution, but worse spectral resolution (e.g. XMM-Newton/EPIC). Let $Y$ be the data from the other instrument, with worse spatial resolution and better spectral resolution (e.g. XRISM/Resolve). When available, $\bar{{X}}$ and $\bar{{Y}}$ are the ground truths at the respective instruments’ resolutions.
In this section, we consider the problem when both HS images have the same size, i.e. $X$ and $Y$ +$\in\mathbb{R}^{M\times N\times L}_{+}$, and thus so does $Z$ (for the case where $X$ and $Y$ have different sizes, see section 2.2). Our forward model (seen in Fig. 2 ) must include each instrument’s spatial degradation, which is modelled by their PSFs, as well as the spectral degradation, modelled by their spectral responses. We must also account for the effective area of the instrument, and the exposure time. For the former, we assume an effective area with no vignetting, i.e. the effective area does not vary spatially. [^8] Accounting for vignetting would be easy, however, since the effective area is simply multiplied to the model term by term. Thus whether we consider the effective area as a spatially uniform cube or a vignetted cube makes no difference to the algorithmic complexity. We call $A_{X}$ and $A_{Y}$ the effective area cubes, which have the same shape as $X$ and $Y$. $t_{X}$ and $t_{Y}$ are scalars, the exposure times of each data set. The instruments’ spectral responses are $\textbf{s}_{X}$ and $\textbf{s}_{Y}$, and their PSFs are $B_{X}$ and $B_{Y}$.
We define $\mathcal{B}_{X}$ and $\mathcal{B}_{Y}$ as the operators which, given a cube $Z$, apply a 2D convolution on its spatial slices with the PSF $B_{X}$ or $B_{Y}$. Similarly, we define as $\mathcal{S}_{X}$ and $\mathcal{S}_{Y}$ the operators which convolve each spectra of $Z$ with $\textbf{s}_{X}$ or $\textbf{s}_{Y}$. Formally, if $Z_{k}$ is the 2D slice of $Z$ at the $k$th spectral channel, and $Z_{ij}$ is the spectra of $Z$ at pixel $(i,j)$:
$\Big{(}\mathcal{B}_{Y}Z\Big{)}_{k}=B_{Y}\circledast Z_{k}$ (3)
@@
The aim is to reach, as best as possible, a fused cube $Z$ that has the spatial resolution of $X$ and the spectral resolution of $Y$. The general problem is:
@@
and the data-fidelity metric $\mathcal{L}_{P}$ is the negative Poisson log-likelihood, $\varphi(Z)$ is the regularization term. We impose $Z\geq 0$ as there is no physical interpretation for a negative spectra. The likelihood terms promote faithfulness between our model and the data at comparable resolutions.
It is worth noting that in order to fully invert the forward model, we would need to compare $X$ with $\mathcal{B}_{X}(t_{X}Z\odot A_{X})\mathcal{S}_{X}$, and likewise for $Y$. However, we are not interested in retrieving a cube $Z$ with infinite resolution, but rather wish $Z$ to have the spatial resolution of $X$, and the spectral resolution $Y$. The problem expressed in Equations 5 and 6 will promote such a solution, since no spatial degradation is applied to go from $Z$ to $Z_{X}$, and likewise spectrally for $Z_{Y}$ with $Y$.
More explicitly, we obtain:
@@ 2.2 Rebinning Operator
In most cases, two different hyperspectral instruments will not produce data of the same voxel size. Rather, we will have $X\in\mathbb{R}_{+}^{M\times N\times l}$, and $Y\in\mathbb{R}_{+}^{m\times n\times L}$, where $L>l$, $M>m$, and $N>n$. In other words, $X$has more pixels than $Y$, and $Y$ has more spectral channels than $X$. Our result $Z$ should be of size $(M\times N\times L)$. In order to compare $Z$ to the data $X$ and $Y$ in the data fidelity terms of equation 5, we thus need to change the size of $Z$ to that of $X$ and $Y$ respectively. This is done by applying a rebinning operator.
For the sake of simplicity, and to keep this operator differentiable, we assume that $M$ is a multiple of $m$, $N$ of $n$, and $L$ of $l$. Let us call the multiplying factors $(\nu_{M},\nu_{N},\nu_{L})$, such that $M=\nu_{M}m$, and likewise for the other dimensions. We need three rebinning operators: $\mathcal{R}_{M}$, $\mathcal{R}_{N}$, and $\mathcal{R}_{L}$.
@@
For instance, for $\mathcal{R}_{M}$, a single voxel $(s,j,k)$ of the rebinned tensor is:
$\big{(}\mathcal{R}_{M}\textbf{Z}\big{)}(s,j,k)=\sum_{p=s*\nu_{M}}^{p=(s+1)*\nu_{M}-1}\textbf{Z}(p,j,k)$ (8)
Such an operator can be neatly written as a matrix product between each slice of the tensor $Z$ and a block diagonal matrix filled only with zeroes and ones. For simplicity, we’ll denote as $\mathcal{R}_{MN}$ the spatial rebinning operator (which applies $\mathcal{R}_{M}$ on the lines, and $\mathcal{R}_{N}$ on the columns).
Now, to include rebinning in our optimization problem posed in equation 5, $Z_{X}$ and $Z_{Y}$ become:
@@
and plugging this in the Poisson likelihood:
@@ 2.3 Regularization
In X-ray astrophysics, our main issue will be the low signal-to-noise ratio, as HS images are dominated by Poisson noise. Thus, we are looking for regularization terms that will be efficient denoisers for Poisson statistics, and which will be appropriate for astrophysical images (often displaying diffuse and/or isotropic features) and X-ray spectra (which can present strong spectral variability).
Our first consideration was sparsity under a 2D-1D starlet transform. Then, inspired by Guilloteau et al. (2022)’s work for JWST, we implemented a low rank approximation with Sobolev spatial regularization. Finally, we wished to test a compromise between the two, with a low rank approximation and a 2D spatial wavelet sparsity regularization. The following sections describe these three options.
A 2D-1D Wavelet Sparsity
A classic regularisation method is to choose a term that promotes the sparsity of $Z$ under a well-chosen orthogonal transform — that is to say, that in some domain, $Z$ can be represented by only a small amount of non-zero coefficients. Wavelets are useful to capture multi-scale features, and in particular, the starlet transform is appropriate for astrophysical data due to its isotropic structure (Starck & Murtagh, 1994).
More specifically, we chose the 2D-1D Starlet transform, which we define as applying the two-dimensional Starlet transform channel-wise, then the one-dimensional Starlet transform pixel-wise. Thus, in this case $\varphi(Z)$ is the $l_{1}$ norm (the minimization of which promotes sparsity, see Bach et al., 2011) of the 2D-1D Starlet transform coefficients of $Z$. For $S$ wavelet scales, we define the 2D-1D starlet transform as $\mathcal{W}_{S}^{2D1D}$, and the $l_{1}$ norm regularization term is:
@@
B Low rank and Sobolev
This regularization option was inspired by Guilloteau et al. (2020)’s work for JWST, though it was implemented from scratch so it could be adapted for HS-HS fusion and the specifics of X-ray astronomy. The spectral regularization is a low-rank approximation, which corresponds to decomposing $Z=WV,$ where $V\in\mathbb{R}^{L\times r}$ spans the subspace of rank $r$ to project $Z$ onto, and $W\in\mathbb{R}^{r\times M\times N}$ is the tensor of representation coefficients in that subspace. $V$ is found by performing principal component analysis on $Y$ (the hyperspectral image with the most spectral information) and keeping only the highest singular values. So now $Z_{X}$ and $Z_{Y}$ from Equation 9 become variables of $W$:
@@ B Low rank and Sobolev
And can be plugged into Equation 10, now seeking to minimize $W$. For spatial regularization, the Sobolev regularization (Karl, 2005) implies adding the following term to the cost function:
@@
where $\mathcal{D}$ is a first order 2-D finite difference operator. This term will thus aim to minimize the differences between neighbouring pixels on the representation coefficients of the subspace spanned by $V$.
Both the Sobolev regularization and the low-rankedness will result in a faster algorithm, due to the dimension reduction and simple expression of the $\mathcal{D}$ operator. However, Sobolev does not capture multi-scale spatial features. Further, since $V$ is extracted from $Y$, it may reproduce features found in $Y$ too strongly. This is fine (and even desirable) in $MS/HS$ fusion, where the MS image does not bring much spectral information, but in $HS/HS$ fusion, $X$ contains spectral information, in particular because it displays much less spatial mixing than $Y$. The assumption that all of $Z$’s spectra can be decomposed as linear combinations of $Y$ is not necessarily true, and forcing a low-rank approximation in this case will bring on spectral distortions. If this assumption is sound, however, low-rankedness will be an undeniable asset.
C Low rank and 2D Wavelet Sparsity
Since Sobolev regularization does not capture multi-scale information the way wavelets do, we considered the option to keep a low-rank approximation as described in the previous section, then to use a starlet sparsity term for the spatial regularization:
@@ 2.4 Algorithm Description
A Proximal Gradient Descent
If we consider the problem posed in Equations 7 or 10, and replace the regularization term by one of the three options we just considered, we notice that the cost function is convex, and can be separated in a differentiable and non-differentiable part. In such cases, proximal gradient methods have been shown to be very effective (Parikh & Boyd, 2014).
The proximal gradient descent is an iterative method where, for each iteration, we first undergo a gradient descent step on the differentiable part of the cost function, then we apply a proximal operator (prox), the purpose of which is to find a solution close to the result obtained by gradient descent while minimizing the function’s non-differentiable part. Formally, if the result obtained by the gradient descent step is $x$, and the non-differentiable part of the function is $\varphi$, the operator is:
@@ A Proximal Gradient Descent
where $\mu$ is the parameter that controls the strength of regularization. For the non-negativity constraint, the prox is simply the operator which puts all negative values to 0.
The gradient of the differentiable part in Equation 5 is:
@@
$=\frac{\partial(Z_{X})}{\partial{Z}}\nabla_{Z_{X}}\Big{(}\mathcal{L}_{P}(X\;|\;Z_{X})\Big{)}+\frac{\partial(Z_{Y})}{\partial{Z}}\nabla_{Z_{Y}}\Big{(}\mathcal{L}_{P}(Y\;|\;Z_{Y})\Big{)}$ (17)
In the case with rebinning, we have:
@@
where $\dagger$ denotes the conjugate transpose of the operators, in the sense of, if we were to vectorize the spatial dimensions of our data and write the operator $\mathcal{S}$ as a product with the 2D matrice $S$, $\mathcal{S}^{\dagger}$ would apply the product with matrix $S^{\dagger}$. In the case with a low rank approximation, where the gradient descent is done on $W=ZV^{T}$:
@@
In the case of Sobolev regularization, $\varphi_{Sobolev}(W)=\mu||\mathcal{D}(W)||^{2}_{F}$ is differentiable, so a term is added to the gradient:
@@
Like the other operators, $\mathcal{D}$ can be interpreted as a product with the 2D matrix $D$, since the operator $\mathcal{D}$ is in fact a convolution with a kernel of $[1,-1]$ on the lines, followed by a convolution with a kernel of $[1,-1]^{T}$ on the columns). $\mathcal{D}^{T}$ applies the product with $D^{T}$.
In the code, the operators $\mathcal{R}$ and $\mathcal{D}$ are written as matrix products. The operators $\mathcal{B}_{Y}$ and $\mathcal{S}_{X}$, however, can be dealt with more efficiently by taking advantage of the diagonalization of the convolution operation in the Fourier domain. To do so, we pad the kernels and $Z$ with zeros so they’re the same size (dimension of $Z$ + dimension of the kernel in question). Then, we take their Fourier transform, and convolution becomes a Hadamar product in the Fourier space. The Fourier transform of $Z$ will need to be calculated at each iteration, but the transform of the kernels need only be calculated once.
@@ 21      $t\leftarrow t+1$
B Hyperparameters
For the choice of hyperparameters (namely, the regularizing factor $\mu$, and the gradient step $\alpha$, and the rank $r$), we need to take a heuristic approach. Indeed, the Poisson likelihood is not globally Lipschitz-continuous, so we cannot analytically calculate an optimal gradient step, and there is no a priori best choice for the regularization parameters either. Further, we found that one set of hyperparameters that worked best for one toy model was not necessarily ideal for another model. The photon count, level of spectral variability, shape of the effective area, and more, are all factors that will impact the choice of best parameters. It would be very costly to compute every single combination of hyperparameters to find the best suited for every single model, but below, we describe some guidelines to help pick suitable parameters, if not the exact optimal value.
The gradient step should be taken as high as possible, for a faster rate of convergence, but if it is taken too high, the likelihood with diverge dramatically. This is easily diagnosed, seen in huge jumps in the cost function, when it should decrease monotonously. This usually happens in early iterations, so we checked the value of the likelihood periodically at the start (first 1000 iterations, taking less than an hour), then stopped if it became pathological. If it was, we then lowered the gradient step and tried again with that parameter.
For the $\mu$ parameter of the Sobolev regularization, for the Gaussian toy model, we tested values between 0.01 and 0.1, and found the best value to be 0.03. This value turned out well suited for the other toy models as well. A too high value of this parameter is detectable by an oversmoothing of spatial features, while a too low value will be recognized in the presence of visible noise on spatial slices.
For the rank $r$, we tested values between 5 and 50, and selected 10. The rank has a similar impact as $\mu$, but on the spectral features. A lower rank might oversmooth and lead to a biased result, while a higher rank might not be enough of a denoiser. On top of that, the rank will impact the algorithm’s computational complexity, with a lower rank implying less dimensions and thus a faster algorihtm.

### v2

@@ 3 Methodology
In this section, we present the methodology of our fusion algorithm for X-ray astrophysics. We present the general problem to be solved in section 3.1. Section 3.2 presents the rebinning operator needed when the data sets have different voxel sizes. Section 3.3 describes our choices of potential regularization terms. Section 3.4 presents our algorithmic strategy to solve the inverse problem.
@@ 3.1 Forward model
Let $X$ be the hyperspectral data from the instrument with higher spatial resolution but lower spectral resolution (e.g., XMM-Newton/EPIC). Let $Y$ be the data from the other instrument, with lower spatial resolution and higher spectral resolution (e.g., XRISM/Resolve). When available, $\bar{{X}}$ and $\bar{{Y}}$ are the ground truths at the resolutions of the respective instruments.
In this section, we consider the problem where both HS images are of the same size, that is, $X$ and $Y$ +$\in\mathbb{R}^{M\times N\times L}_{+}$, and therefore so does $Z$ (see section 3.2 for the case where $X$ and $Y$ have different sizes). Our forward model (seen in Fig. 2) must include the spatial degradation of each instrument, which is modeled by their PSFs, as well as the spectral degradation, which is modeled by their spectral responses. We must also account for the effective area of the instrument, and the exposure time. For the former, we assume an effective area with no vignetting; that is, the effective area does not vary spatially. [^8] However, accounting for vignetting would be easy, since the effective area is simply multiplied to the model term by term. Thus whether we consider the effective area as a spatially uniform cube or a vignetted cube has no influence on the algorithmic complexity. We use $A_{X}$ and $A_{Y}$ to denote the effective area cubes, which have the same shape as $X$ and $Y$. $t_{X}$ and $t_{Y}$ are scalars, the exposure times of each data set. The spectral responses of the instruments are $\textbf{s}_{X}$ and $\textbf{s}_{Y}$, and their PSFs are $B_{X}$ and $B_{Y}$.
We define $\mathcal{B}_{X}$ and $\mathcal{B}_{Y}$ as the operators which, given a cube $Z$, apply a 2D convolution on its spatial slices with the PSF $B_{X}$ or $B_{Y}$. Similarly, we use $\mathcal{S}_{X}$ and $\mathcal{S}_{Y}$ to denote the operators that convolve each spectra of $Z$ with $\textbf{s}_{X}$ or $\textbf{s}_{Y}$, respectively. Formally, if $Z_{k}$ is the 2D slice of $Z$ at the $k$th spectral channel, and $Z_{ij}$ is the spectra of $Z$ at pixel $(i,j)$:
$\Big{(}\mathcal{B}_{Y}Z\Big{)}_{k}=B_{Y}\circledast Z_{k},$ (3)
@@
The aim is to reach, as closely as possible given the noisy data, a fused cube $Z$ that has the spatial resolution of $X$ and the spectral resolution of $Y$. The general problem is
@@
where $\odot$ is the term-by-term product, the data-fidelity metric $\mathcal{L}_{P}$ is the negative Poisson log-likelihood, and $\varphi(Z)$ is the regularization term. We impose $Z\geq 0,$ as there is no physical interpretation for a negative spectra. The likelihood terms promote faithfulness between our model and the data at comparable resolutions.
It is worth noting that in order to fully invert the forward model, we would need to compare $X$ with $\mathcal{B}_{X}(t_{X}Z\odot A_{X})\mathcal{S}_{X}$, and likewise for $Y$. However, our aim is not to retrieve a cube $Z$ with infinite resolution, but rather our intention is for $Z$ to have the spatial resolution of $X$, and the spectral resolution $Y$. The problem expressed in Equations 5 and 6 will promote such a solution, since no spatial degradation is applied to go from $Z$ to $Z_{X}$, and likewise spectrally for $Z_{Y}$ with $Y$.
More explicitly, we obtain
@@ 3.2 Rebinning operator
In most cases, two different hyperspectral instruments will not produce data of the same voxel size. Rather, we will have $X\in\mathbb{R}_{+}^{M\times N\times l}$ and $Y\in\mathbb{R}_{+}^{m\times n\times L}$, where $L>l$, $M>m$, and $N>n$. In other words, $X$ has more pixels than $Y$, and $Y$ has more spectral channels than $X$. Our result $Z$ should be of size $(M\times N\times L)$. In order to compare $Z$ to the data $X$ and $Y$ in the data fidelity terms of equation 5, we thus need to change the size of $Z$ to that of $X$ and $Y,$ respectively. This is done by applying a rebinning operator.
For the sake of simplicity, and to keep this operator differentiable, we assume that $M$ is a multiple of $m$, $N$ is a multiple of $n$, and $L$ is a multiple of $l$. Let us denote the multiplying factors $(\nu_{M},\nu_{N},\nu_{L})$, such that $M=\nu_{M}m$, and likewise for the other dimensions. We need three rebinning operators: $\mathcal{R}_{M}$, $\mathcal{R}_{N}$, and $\mathcal{R}_{L}$.
@@
For instance, for $\mathcal{R}_{M}$, a single voxel $(s,j,k)$ of the rebinned tensor is
$\big{(}\mathcal{R}_{M}\textbf{Z}\big{)}(s,j,k)=\sum_{p=s*\nu_{M}}^{p=(s+1)*\nu_{M}-1}\textbf{Z}(p,j,k).$ (8)
Such an operator can be neatly written as a matrix product between each slice of the tensor $Z$ and a block diagonal matrix filled only with zeroes and ones. For simplicity, we use $\mathcal{R}_{MN}$ to denote the spatial rebinning operator (which applies $\mathcal{R}_{M}$ on the lines, and $\mathcal{R}_{N}$ on the columns).
Now, to include rebinning in our optimization problem posed in equation 5, $Z_{X}$ and $Z_{Y}$ become
@@
and plugging this into the Poisson likelihood, we obtain
@@ 3.3 Regularization
In X-ray astrophysics, low signal-to-noise ratio presents a major problem, as HS images are dominated by Poisson noise. Thus, we are looking for regularization terms that will be efficient denoisers for Poisson statistics, and that will be appropriate for astrophysical images (often displaying diffuse and/or isotropic features) and X-ray spectra (which can present strong spectral variability).
Our first consideration was sparsity under a 2D-1D starlet transform. Then, inspired by the work of Guilloteau et al. (2022) regarding JWST, we implemented a low-rank approximation with Sobolev spatial regularization. Finally, we wished to test a compromise between the two, with a low-rank approximation and a 2D spatial wavelet sparsity regularization. The following sections describe these three options.
3.3.1 2D-1D wavelet sparsity
A classic regularization method is to choose a term that promotes the sparsity of $Z$ under a well-chosen orthogonal transform — that is to say, that in some domain, $Z$ can be represented by only a small number of non-zero coefficients. Wavelets are useful for capturing multi-scale features. In particular, the isotropic undecimated wavelet transform, dubbed the starlet transform, is appropriate for astrophysical data due to its isotropic structure (Starck & Murtagh, 1994). The data are convolved by $B_{3}$-spline filters, which produce a set of increasingly coarse resolution versions of the data, and the $S$ wavelet scales are defined as the difference between two subsequent resolutions. The coarsest resolution is dubbed the coarse scale, and no assumption of sparsity is placed upon it. This allows the regularization to represent data with a smooth continuum (the coarse scale) and sharp features (the non-zero elements of the fine scales).
More specifically, we chose the 2D-1D starlet transform, which we define as applying the 2D starlet transform channel-wise, then the 1D starlet transform pixel-wise. Thus, in this case, $\varphi(Z)$ is the $l_{1}$ norm (the minimization of which promotes sparsity; see Bach et al., 2011) of the 2D-1D starlet transform coefficients of $Z$. For $S$ wavelet scales, we define the 2D-1D starlet transform as $\mathcal{W}_{S}^{2D1D}$, and the $l_{1}$ norm regularization term is
@@
3.3.2 Low rank and Sobolev
This regularization option was inspired by work regarding JWST carried out by Guilloteau et al. (2020), though we designed an implementation adapted for HS-HS fusion and the specifics of X-ray astronomy. The spectral regularization is a low-rank approximation, which corresponds to decomposing $Z=WV,$ where $V\in\mathbb{R}^{L\times r}$ spans the subspace of rank $r$ onto which $Z$ is projected, and $W\in\mathbb{R}^{r\times M\times N}$ is the tensor of representation coefficients in that subspace. $V$ is found by performing principal component analysis on $Y$ (the hyperspectral image with the most spectral information) and keeping only the highest singular values. Therefore, now $Z_{X}$ and $Z_{Y}$ from Equation 9 become variables of $W$, where
@@ 3.3.2 Low rank and Sobolev
and these can be plugged into Equation 10, now seeking to minimize $W$. For spatial regularization, the Sobolev regularization (Karl, 2005) implies adding the following term to the cost function:
@@
where $\mathcal{D}$ is a first-order 2D finite difference operator. This term will act to minimize the differences between neighboring pixels on the representation coefficients of the subspace spanned by $V$.
Both the Sobolev regularization and the low-rankedness will result in a faster algorithm, due to the dimension reduction and simple expression of the $\mathcal{D}$ operator. However, Sobolev does not capture multi-scale spatial features. Further, as $V$ is extracted from $Y$, it may reproduce features found in $Y$ too strongly. This is acceptable (and even desirable) in $MS/HS$ fusion, where the MS image does not bring a significant amount of spectral information, but in $HS/HS$ fusion, $X$ contains spectral information, in particular because it displays much less spatial mixing than $Y$. The assumption that all spectra in $Z$ can be decomposed as linear combinations of $Y$ is not necessarily true, and forcing a low-rank approximation in this case will bring on spectral distortions. However, if this assumption is valid, low-rankedness will be an undeniable asset.
3.3.3 Low rank and 2D wavelet sparsity
As Sobolev regularization does not capture multi-scale information in the same way wavelets do, we considered the option to keep a low-rank approximation as described in the previous section, and then to use a starlet sparsity term for the spatial regularization:
@@ 3.4 Algorithm description
3.4.1 Proximal gradient descent
If we consider the problem posed in Equations 7 or 10, and replace the regularization term by one of the three options we consider above, we note that the cost function is convex, and can be separated into two parts, one differentiable and one non-differentiable. In such cases, proximal gradient methods have been shown to be very effective (Parikh & Boyd, 2014).
The proximal gradient descent is an iterative method where, for each iteration, we first undergo a gradient descent step on the differentiable part of the cost function before applying a proximal operator (prox), the purpose of which is to find a solution close to the result obtained by gradient descent while minimizing the function’s non-differentiable part. Formally, if the result obtained by the gradient descent step is $x$, and the non-differentiable part of the function is $\varphi$, the operator is:
@@ 3.4.1 Proximal gradient descent
where $\mu$ is the parameter that controls the strength of regularization. For the non-negativity constraint, the prox is simply the operator that forces all negative values to 0.
The gradient of the differentiable part in Equation 5 is
@@
$=\frac{\partial(Z_{X})}{\partial{Z}}\nabla_{Z_{X}}\Big{(}\mathcal{L}_{P}(X\;|\;Z_{X})\Big{)}+\frac{\partial(Z_{Y})}{\partial{Z}}\nabla_{Z_{Y}}\Big{(}\mathcal{L}_{P}(Y\;|\;Z_{Y})\Big{)}.$ (17)
In the case with rebinning, we have
@@
where $\dagger$ denotes the conjugate transpose of the operators, in the sense that, if we were to vectorize the spatial dimensions of our data and write the operator $\mathcal{S}$ as a product with the 2D matrice $S$, $\mathcal{S}^{\dagger}$ would apply the product with matrix $S^{\dagger}$. In the case with a low-rank approximation, where the gradient descent is done on $W=ZV^{T}$:
@@
In the case of Sobolev regularization, $\varphi_{Sobolev}(W)=\mu||\mathcal{D}(W)||^{2}_{F}$ is differentiable, and so a term is added to the gradient:
@@
As is the case for the other operators, $\mathcal{D}$ can be interpreted as a product with the 2D matrix $D$, since the operator $\mathcal{D}$ is a convolution with a kernel of $[1,-1]$ on the lines followed by a convolution with a kernel of $[1,-1]^{T}$ on the columns. $\mathcal{D}^{T}$ applies the product with $D^{T}$.
In the code, the operators $\mathcal{R}$ and $\mathcal{D}$ are written as matrix products. The operators $\mathcal{B}_{Y}$ and $\mathcal{S}_{X}$, however, can be dealt with more efficiently by taking advantage of the diagonalization of the convolution operation in the Fourier domain. To do so, we pad the kernels and $Z$ with zeros so that they are the same size (dimension of $Z$ + dimension of the kernel in question). We then take their Fourier transform, and convolution becomes a Hadamar product in the Fourier space. The Fourier transform of $Z$ will need to be calculated at each iteration, but the transform of the kernels need only be calculated once.
@@ 21      $t\leftarrow t+1$
3.4.2 Hyperparameters
For the choice of hyperparameters (namely the regularizing factor $\mu$, the gradient step $\alpha$, and the rank $r$), we need to take a heuristic approach. Indeed, the Poisson likelihood is not globally Lipschitz-continuous, and so we cannot analytically calculate an optimal gradient step, and there is no a priori best choice for the regularization parameters either. Further, we find that one set of hyperparameters that works best for one toy model is not necessarily ideal for another model. The photon count, level of spectral variability, shape of the effective area, and more are all factors that will impact the choice of best parameters. It would be very costly to compute every single combination of hyperparameters to find the best suited for every single model, but below, we describe some guidelines to help pick suitable parameters, if not the exact optimal value.
The gradient step should be taken as high as possible for a faster rate of convergence, but if it is taken too high, the likelihood will diverge dramatically. This is easily diagnosed, and is seen as huge jumps in the cost function, when it should rather decrease monotonously. This usually happens in early iterations, and so we checked the value of the likelihood periodically at the start (first 1000 iterations, which takes less than an hour), and then stopped if it became pathological. If it this was the case, we then lowered the gradient step and tried again with that parameter. As the Poisson likelihood is convex, there is no risk of falling into a local minimum, provided this step is well-chosen as described.
For the $\mu$ parameter of the Sobolev regularization, for the Gaussian toy model, we tested values of between 0.01 and 0.1, and found the best value to be 0.03. This value turned out to be well suited for the other toy models as well. An overly high value of this parameter is detectable as an over-smoothing of spatial features, while an excessively low value will be recognized in the presence of visible noise on spatial slices.
For the rank $r$, we tested values of between 5 and 50, and selected 10. The rank has a similar impact to $\mu$, but on the spectral features. A lower rank might over-smooth and lead to a biased result, while a higher rank might not be enough of a denoiser. On top of that, the rank will impact the algorithm’s computational complexity, with a lower rank implying fewer dimensions and thus a faster algorithm.